OpenStudy (anonymous):

Find the following products. a. (2x-1)(x+3) b. 3x(5x-11) c. (x-5)(5x-2) d. 100(3x-0.5)

OpenStudy (anonymous):

a=7x-3 b=15x-33

OpenStudy (anonymous):

Thanks. What about c and d ?

OpenStudy (texaschic101):

a. (2x - 1)(x + 3) = 2x^2 + 6x - x - 3 = 2x^2 + 5x - 3 b. 3x(5x - 11) 15x^2 - 33x c. (x - 5)(5x - 2) 5x^2 - 2x - 25x + 10 = 5x^2 - 27x + 10 d. 100 (3x - 0.5) = 300x - 50

OpenStudy (anonymous):

Thank you! (:

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

c=-7x-10 d=300x-50

OpenStudy (anonymous):

Thanks!

OpenStudy (texaschic101):

reneebaby...how are you getting your answers for a and c ??

OpenStudy (texaschic101):

@reneebaby139 ...

OpenStudy (texaschic101):

a and c require the FOIL method to solve http://www.algebrahelp.com/lessons/simplifying/foilmethod/

OpenStudy (anonymous):

I was using foil, it may be wrong though I had algebra last yr and we haven't done it since then. My teacher called it the rainbow method, bc you distribute everything then add/subtract similar things that you can. for (x-5)5x-2), I got 5x-2x-10x-10, then I added like terms to get -7x-10